# Defining implication relation for classical logic

### v11

@@ 1 $\phi\wedge\psi\leftrightarrow\phi$ Premise
2     $\leftrightarrow(\phi\wedge\psi)\vee\psi$ 1: Replacement
@@ 4     $\leftrightarrow(\phi\vee\psi)\wedge\neg\psi$ Distributivity
5     $\leftrightarrow\psi\wedge\neg\psi$ 1: Replacement
@@ 2     $\leftrightarrow\neg\phi\vee\bot\vee\psi$ Identity
3     $\leftrightarrow\neg\phi\vee(\phi\wedge\neg\psi)\vee\psi$ 1: Replacement
@@ 15 $\neg(\phi\wedge\psi)\leftrightarrow(\neg\phi\vee\neg\psi)$ 7,14: Bi-implication introduction
Theorem 3.4.12 (One-to-two and Two-to-one).
$(\chi\to\phi)\wedge(\chi\to\psi)\leftrightarrow(\chi\to\phi\wedge\psi),\;(\phi\to\chi)\wedge(\psi\to\chi)\leftrightarrow(\phi\vee\psi\to\chi).$
@@ 9     $\phi\wedge\psi\to\phi$ Conjunction elimination
10     $\chi\to\phi$ 8–9: Transitivity
11     $\chi\to\psi$ 8–10: similarly
12     $(\chi\to\phi)\wedge(\chi\to\psi)$ 10,11: Conjunction introduction
13 $(\chi\to\phi\wedge\psi)\to(\chi\to\phi)\wedge(\chi\to\psi)$ 8–12: Conditional proof
14 $(\chi\to\phi)\wedge(\chi\to\psi)\leftrightarrow(\chi\to\phi\wedge\psi)$ 7,13: Bi-implication introduction
@@ 13 $(\chi\to\phi\wedge\psi)\to(\chi\to\phi)\wedge(\chi\to\psi)$ 8–12: Conditional proof
Theorem 3.4.13 (Strengthening).
$(\phi\to\psi)\to(\phi\to\phi\wedge\psi),\;(\phi\to\psi)\to(\phi\vee\psi\to\psi).$
@@ 3 $\phi\to\phi\wedge\psi$ 1,2: One-to-two
Theorem 3.4.14 (Weakening).
$(\phi\to\psi)\to(\phi\wedge\chi\to\psi),\;(\phi\to\psi)\to(\phi\to\psi\vee\chi).$
@@ 3 $\phi\wedge\chi\to\psi$ 2,1: Transitivity
Theorem 3.4.15 (Compatibility of implication with conjunction and disjunction).
$(\phi\to\psi)\to(\phi\wedge\chi\to\psi\wedge\chi),\;(\phi\to\psi)\to(\phi\vee\chi\to\psi\vee\chi).$
@@ 4 $\phi\wedge\chi\to\psi\wedge\chi$ 2,3: One-to-two
Theorem 3.4.16 (Monotonicity of conjunction and disjunction w.r.t. implication).
$(\phi_{1}\to\psi_{1})\wedge(\phi_{2}\to\psi_{2})\to(\phi_{1}\wedge\phi_{2}\to\psi_{1}\wedge\psi_{2}),$
$(\phi_{1}\to\psi_{1})\wedge(\phi_{2}\to\psi_{2})\to(\phi_{1}\vee\phi_{2}\to\psi_{1}\vee\psi_{2}).$
@@ 5 $\phi_{1}\wedge\phi_{2}\to\psi_{1}\wedge\psi_{2}$ 3,4: One-to-two
Corollary 3.4.1 (Constructive dilemma).
$(\phi_{1}\vee\phi_{2})\wedge(\phi_{1}\to\psi_{1})\wedge(\phi_{2}\to\psi_{2})\to\psi_{1}\vee\psi_{2}.$
Proof.
1 $\phi_{1}\vee\phi_{2}$ Premise
2 $\phi_{1}\to\psi_{1}$ Premise
3 $\phi_{2}\to\psi_{2}$ Premise
5 $\psi_{1}\vee\psi_{2}$ 1,4: Modus penens
@@ 1     $\leftrightarrow(\phi\wedge\neg\psi\leftrightarrow\bot)$ Definition of implication
Theorem 3.4.19 (Bi-implication elimination).
$(\phi\leftrightarrow\psi)\to(\phi\to\psi).$
@@ 3 $\phi\to\psi$ 1,2: Replacement
Theorem 3.4.20 (Bi-implication by implications).
$(\phi\leftrightarrow\psi)\leftrightarrow(\phi\to\psi)\wedge(\psi\to\phi).$
@@ 2 $\phi\leftrightarrow\psi$ 1: Bi-implication introduction
Theorem 3.4.21 (Bi-implication of negations).
$(\phi\leftrightarrow\psi)\leftrightarrow(\neg\phi\leftrightarrow\neg\psi).$
@@ 1     $\leftrightarrow(\phi\to\psi)\wedge(\psi\to\phi)$ Bi-implication by implications
Bounds of implication
Theorem 3.4.22 (Implication-to-Disjunction / Maximum of implication).
$(\phi\to\psi)\to\neg\phi\vee\psi.$
Proof.
1 $\phi\to\psi$ Premise
3 $\top$ Top is a theorem
@@ 5 $\chi\to(\phi\to\psi)$ 2,4: Replacement $[\phi\wedge\psi\mapsto\phi]$
Note: Generalized ECQ $\psi\to(\phi\to\psi)$ does not hold for IRL (cf. Subsection 5.3).
@@
1 $\psi\leftrightarrow\top$ Premise
2 ϕitalic-$\phi\to\top$ Dual of ECQ
3 $\phi\to\psi$ 1,2: Replacement $[\top\mapsto\psi]$
@@ 4 $\phi\wedge\psi\to\chi$ 2–3: Transitivity
Exportation (converse of importation) $(\phi\wedge\psi\to\chi)\to(\phi\to(\psi\to\chi))$ does not hold for IRL except for some special forms.
@@ 5 $(\psi\to\phi)\to(\phi\leftrightarrow\psi)$ 3,4: Transitivity
Theorem 3.4.27 (Exported transitivity of implication).
@@ 1 $\phi\to\psi$ Premise
$(\psi\to\chi)$
2     $\to(\phi\to\psi)$ 1: Generalized ECQ of implication
3     $\to(\phi\to\psi)\wedge(\psi\to\chi)$ 2: Strengthening
4     $\to(\phi\to\chi)$ Transitivity of implication
5 $(\psi\to\chi)\to(\phi\to\chi)$ 2–4: Transitivity of implication
The second is obtained via contraposition.
@@ 1 $\chi\to\phi$ Premise
$(\chi\to\psi)$
2     $\to(\chi\to\phi)$ 1: Generalized ECQ of implication
3     $\to(\chi\to\phi)\wedge(\chi\to\psi)$ 2: Strengthening
4     $\to(\chi\to\phi\wedge\psi)$ One-to-two
5 $(\chi\to\psi)\to(\chi\to\phi\wedge\psi))$ 2–4: Transitivity
@@ 4     $\leftrightarrow(\phi\to\bot)\wedge(\bot\to\phi)$ Bi-implication by implication
Theorem 3.4.31 (Maximum of univariate implication).
@@
They follow from Theorem 3.4.22 by uniform substitution, and properties of top and bottom. ∎
@@
It follows from Theorem 3.4.31. ∎
@@
Theorem 3.4.32 (Disjunctive syllogism).
$(\phi\vee\psi)\wedge\neg\phi\to\psi.$
@@
1 $(\phi\vee\psi)\wedge\neg\phi$ Premise
2 $(\phi\wedge\neg\phi)\vee(\psi\wedge\neg\phi)$ 1: Distributivity
3 $\bot\vee(\psi\wedge\neg\phi)$ 2: LNC
4 $\psi\wedge\neg\phi$ 3: Identity element
5 $\psi$ 4: Conjunction elimination
Theorem 3.4.33 (Proof by exhaustion).
$(\phi\vee\psi)\wedge(\phi\to\chi)\wedge(\psi\to\chi)\to\chi.$
@@ 1 $\phi\vee\psi$ Premise
2 $(\phi\to\chi)\wedge(\psi\to\chi)$ Premise
3 $\phi\vee\psi\to\chi$ 2: Two-to-one
4 $\chi$ 1,3: Modus ponens
Theorem 3.4.34 (RAA (reductio ad absurdum)).
$(\phi\to\psi)\wedge(\phi\to\neg\psi)\to\neg\phi.$
@@ 4 $\chi$ 1,3: Modus ponens
1 $(\phi\to\psi)\wedge(\phi\to\neg\psi)$ Premise
2 $\phi\to\psi\wedge\neg\psi$ 1: One-to-two
3 ϕitalic-$\phi\to\bot$ LNC,2: Replacement
4 $\neg\phi$ 3: Maximum of univariate implication
All of $(\phi\to\psi)\wedge(\phi\to\neg\psi)\to\neg\phi$, $(\phi\to\neg\phi)\to\neg\phi$ and $(\phi\to\bot)\to\neg\phi$ are forms of RAA, while ($(\neg\phi\to\psi)\wedge(\neg\phi\to\neg\psi)\to\phi$, ($(\neg\phi\to\phi)\to\phi$ (consequentia mirabilis) and ($(\neg\phi\to\bot)\to\phi$ are forms of proof by contradiction.
Lemma 3.4.2.
$(\chi\wedge\phi\to\psi)\wedge(\neg\chi\wedge\phi\to\psi)\to(\phi\to\psi).$
@@ 4 $\neg\phi$ 3: Maximum of univariate implication
1 $(\chi\wedge\phi\to\psi)\wedge(\neg\chi\wedge\phi\to\psi)$ Premise
2 $(\chi\wedge\phi)\vee(\neg\chi\wedge\phi)\to\psi$ 1: Two-to-one
3 $(\chi\vee\neg\chi)\wedge\phi\to\psi$ Distributivity,2: Replacement
4 $\phi\to\psi$ Theorem (LEM) like top,3: Replacement
Theorem 3.4.35 (Resolution).
$(\phi\vee\chi)\wedge(\psi\vee\neg\chi)\to\phi\vee\psi.$
@@ 1 $\neg\chi\wedge(\phi\vee\chi)\to\phi$ Disjunctive syllogism
2 $\neg\chi\wedge(\phi\vee\chi)\wedge(\psi\vee\neg\chi)\to\phi$ 1: Weakening
3 $\neg\chi\wedge(\phi\vee\chi)\wedge(\psi\vee\neg\chi)\to\phi\vee\psi$ 2: Strengthening
4 $\chi\wedge(\phi\vee\chi)\wedge(\psi\vee\neg\chi)\to\phi\vee\psi$ 1–3: similarly
5 $(\phi\vee\chi)\wedge(\psi\vee\neg\chi)\to\phi\vee\psi$ 3,4: Lemma 3.4.2
Theorem 3.4.36 (Cut).
$(\phi_{1}\to\chi)\wedge(\phi_{2}\wedge\chi\to\psi)\to(\phi_{1}\wedge\phi_{2}\to\psi).$

### v12

@@ 1 $\phi\wedge\psi\leftrightarrow\phi$ Premise
2     $\leftrightarrow(\phi\wedge\psi)\vee\psi$ 1: Replacement $[\phi\mapsto\phi\wedge\psi]$
@@ 4     $\leftrightarrow(\phi\vee\psi)\wedge\neg\psi$ Distributivity
5     $\leftrightarrow\psi\wedge\neg\psi$ 1: Replacement $[\phi\vee\psi\mapsto\psi]$
@@ 2     $\leftrightarrow\neg\phi\vee\bot\vee\psi$ Identity
3     $\leftrightarrow\neg\phi\vee(\phi\wedge\neg\psi)\vee\psi$ 1: Replacement $[\bot\mapsto\phi\wedge\neg\psi]$
@@ 2 $\phi\to\psi$ 1: Definition of implication
Implication-to-Disjunction
Theorem 3.4.7 (Implication-to-Disjunction / Upper bound of implication).
$(\phi\to\psi)\to\neg\phi\vee\psi.$
Proof.
1 $\phi\to\psi$ Premise
3 $\top$ Top is a theorem
@@ 15 $\neg(\phi\wedge\psi)\leftrightarrow(\neg\phi\vee\neg\psi)$ 7,14: Bi-implication introduction
Rule 3.4.3 (One-to-two and Two-to-one).
$\{\chi\to\phi,\chi\to\psi\}\Leftrightarrow\chi\to\phi\wedge\psi,\;\{\phi\to\chi,\psi\to\chi\}\Leftrightarrow\phi\vee\psi\to\chi.$
@@ 2 $\phi\wedge\psi\to\phi$ Conjunction elimination
3 $\chi\to\phi$ 1–2: Transitivity
4 $\chi\to\psi$ 1–3: Similarly
@@
Rule 3.4.4 (Strengthening).
$\phi\to\psi\Rightarrow\phi\to\phi\wedge\psi,\;\phi\to\psi\Rightarrow\phi\vee\psi\to\psi.$
@@ 3 $\phi\to\phi\wedge\psi$ 1,2: One-to-two
Rule 3.4.5 (Weakening).
$\phi\to\psi\Rightarrow\phi\wedge\chi\to\psi,\;\phi\to\psi\Rightarrow\phi\to\psi\vee\chi.$
@@ 3 $\phi\wedge\chi\to\psi$ 2,1: Transitivity
Rule 3.4.6 (Compatibility of implication with conjunction and disjunction).
$\phi\to\psi\Rightarrow\phi\wedge\chi\to\psi\wedge\chi,\;\phi\to\psi\Rightarrow\phi\vee\chi\to\psi\vee\chi.$
@@ 4 $\phi\wedge\chi\to\psi\wedge\chi$ 2,3: One-to-two
Rule 3.4.7 (Monotonicity of conjunction and disjunction w.r.t. implication).
$\{\phi_{1}\to\psi_{1},\phi_{2}\to\psi_{2}\}\Rightarrow\phi_{1}\wedge\phi_{2}\to\psi_{1}\wedge\psi_{2},$
$\{\phi_{1}\to\psi_{1},\phi_{2}\to\psi_{2}\}\Rightarrow\phi_{1}\vee\phi_{2}\to\psi_{1}\vee\psi_{2}.$
@@ 1     $\leftrightarrow(\phi\wedge\neg\psi\leftrightarrow\bot)$ Definition of implication
4 $(\phi\to\psi)\leftrightarrow(\neg\psi\to\neg\phi)$ 1–3: Transitivity
Rule 3.4.8 (Modus tollens).
$\{\phi\to\psi,\neg\psi\}\Rightarrow\neg\phi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\neg\psi$ Premise
3 $\neg\psi\to\neg\phi$ 1: Contraposition
4 $\neg\phi$ 2,3: Modus ponens
@@
Rule 3.4.9 (Bi-implication elimination).
$\phi\leftrightarrow\psi\Rightarrow\phi\to\psi.$
@@ 3 $\phi\to\psi$ 1,2: Replacement
Rule 3.4.10 (Bi-implication by implications).
$\phi\leftrightarrow\psi\Leftrightarrow(\phi\to\psi)\wedge(\psi\to\phi).$
@@ 2 $\phi\leftrightarrow\psi$ 1: Bi-implication introduction
Rule 3.4.11 (Bi-implication of negation).
$\phi\leftrightarrow\psi\Leftrightarrow\neg\phi\leftrightarrow\neg\psi.$
@@ 5 $\chi\to(\phi\to\psi)$ 2,4: Replacement $[\phi\wedge\psi\mapsto\phi]$
It can also be derived in a meta proof:
1 $\vdash\phi\wedge\psi\to\psi$ Conjunction elimination 2 $\vdash\chi\to(\phi\wedge\psi\to\psi)$ 1: Theorem like top 3     $\phi\to\psi$ Premise 4     $\phi\wedge\psi\leftrightarrow\phi$ 3: Definition of implication 5     $\chi\to(\phi\wedge\psi\to\psi)$ 2 (theorem from an outer layer) 6     $\chi\to(\phi\to\psi)$ 4,5: Replacement $[\phi\wedge\psi\mapsto\phi]$ 7 $\vdash(\phi\to\psi)\to(\chi\to(\phi\to\psi))$ 3–6: Conditional proof
Note that Generalized ECQ of the general form $\psi\to(\phi\to\psi)$ does not hold for IRL (cf. Subsection 5.3).
@@
1 $\top\to\phi$ Premise
2 $\psi\to\top$ Dual of ECQ
3 $\psi\to\phi$ 2,1: Transitivity
@@ 4 $\phi\wedge\psi\to\chi$ 2–3: Transitivity
Exportation (converse of importation) $(\phi\wedge\psi\to\chi)\to(\phi\to(\psi\to\chi))$ does not hold for IRL except for some special cases as follows.
@@ 5 $(\psi\to\phi)\to(\phi\leftrightarrow\psi)$ 3,4: Transitivity
Theorem 3.4.16 (Exported transitivity of implication / Suffixing and Prefixing).
@@ 1 $\phi\to\psi$ Premise
2 $(\psi\to\chi)\to(\phi\to\psi)$ 1: Generalized ECQ of implication
3 $(\psi\to\chi)\to(\phi\to\psi)\wedge(\psi\to\chi)$ 2: Strengthening
4 $(\phi\to\psi)\wedge(\psi\to\chi)\to(\phi\to\chi)$ Transitivity of implication
5 $(\psi\to\chi)\to(\phi\to\chi)$ 3–4: Transitivity of implication
The second is obtained via uniform substitution and contraposition.
@@ 1 $\chi\to\phi$ Premise
2 $(\chi\to\psi)\to(\chi\to\phi)$ 1: Generalized ECQ of implication
3 $(\chi\to\psi)\to(\chi\to\phi)\wedge(\chi\to\psi)$ 2: Strengthening
4 $(\chi\to\phi)\wedge(\chi\to\psi)\to(\chi\to\phi\wedge\psi)$ One-to-two
5 $(\chi\to\psi)\to(\chi\to\phi\wedge\psi))$ 3–4: Transitivity
@@ 4     $\leftrightarrow(\phi\to\bot)\wedge(\bot\to\phi)$ Bi-implication by implication
Theorem 3.4.20 (Upper bound of univariate implication).
@@
They follow from the upper bound of implication $(\phi\to\psi)\to\neg\phi\vee\psi$ (Theorem 3.4.7) by uniform substitution, and properties of top and bottom. ∎
@@
It follows from Theorem 3.4.20 by uniform substitution. ∎
@@
Rule 3.4.12 (Constructive dilemma).
$\{\phi_{1}\vee\phi_{2},\phi_{1}\to\psi_{1},\phi_{2}\to\psi_{2}\}\Rightarrow\psi_{1}\vee\psi_{2}.$
@@
1 $\phi_{1}\vee\phi_{2}$ Premise
2 $\phi_{1}\to\psi_{1}$ Premise
3 $\phi_{2}\to\psi_{2}$ Premise
5 $\psi_{1}\vee\psi_{2}$ 1,4: Modus penens
Rule 3.4.13 (Proof by exhaustion).
$\{\phi\vee\psi,\phi\to\chi,\psi\to\chi\}\Rightarrow\chi.$
@@ 1 $\phi\vee\psi$ Premise
2 $\phi\to\chi$ Premise
3 $\psi\to\chi$ Premise
4 $\phi\vee\psi\to\chi$ 2,3: Two-to-one
5 $\chi$ 1,4: Modus ponens
Rule 3.4.14 (LEM exhaustion).
$\{\phi\to\psi,\neg\phi\to\psi\}\Rightarrow\psi.$
@@ 5 $\chi$ 1,4: Modus ponens
1 $\phi\to\psi$ Premise
2 $\neg\phi\to\psi$ Premise
3 $\psi\vee\neg\phi$ LEM
4 $\psi$ 3,1,2: Proof by exhaustion
Rule 3.4.15 (Reductio ad absurdum (RAA)).
$\{\phi\to\psi,\phi\to\neg\psi\}\to\neg\phi.$
@@ 4 $\psi$ 3,1,2: Proof by exhaustion
1 $\phi\to\psi$ Premise
2 $\phi\to\neg\psi$ Premise
3 $\neg\psi\to\neg\phi$ 1: Contraposition
4 $\psi\to\neg\phi$ 2: Contraposition, Double negation
5 $\neg\phi$ 3,4: LEM exhaustion
The rules $\{\phi\to\psi,\phi\to\neg\psi\}\Rightarrow\neg\phi$, $\phi\to\neg\phi\Rightarrow\neg\phi$ and $\phi\to\bot\Rightarrow\neg\phi$ are all forms of RAA, while the rules $\{\neg\phi\to\psi,\neg\phi\to\neg\psi\}\Rightarrow\phi$, $\neg\phi\to\phi\Rightarrow\phi$ (consequentia mirabilis) and $\neg\phi\to\bot\Rightarrow\phi$ are forms of proof by contradiction.
Rule 3.4.16 (Disjunctive syllogism).
$\{\phi\vee\psi,\neg\phi\}\Rightarrow\psi.$
Proof.
1 $\phi\vee\psi$ Premise
2 $\neg\phi$ Premise
3 $(\phi\vee\psi)\wedge\neg\phi$ 1,2: Conjunction introduction
4 $(\phi\wedge\neg\phi)\vee(\psi\wedge\neg\phi)$ 3: Distributivity
5 $\phi\wedge\neg\phi\to\psi$ ECQ
6 $\psi\wedge\neg\phi\to\psi$ Conjunction elimination
7 $\psi$ 4,5,6: Proof by exhaustion
Rule 3.4.17 (Implication and inconsistency).
$\phi\to\psi\Leftrightarrow\phi\wedge\neg\psi\to\bot.$
Proof.
$(\phi\to\psi)$
2     $\leftrightarrow(\phi\wedge\neg\psi\to\bot)$ Univariate implication expressions
3 $(\phi\to\psi)\leftrightarrow(\phi\wedge\neg\psi\to\bot)$ 1,2: Transitivity
Rule 3.4.18 (LEM exhaustion for implication).
$\{\chi\wedge\phi\to\psi,\neg\chi\wedge\phi\to\psi\}\Rightarrow\phi\to\psi.$
Proof.
1 $\chi\wedge\phi\to\psi$ Premise
2 $\neg\chi\wedge\phi\to\psi$ Premise
3 $(\chi\wedge\phi)\vee(\neg\chi\wedge\phi)\to\psi$ 1,2: Two-to-one
$(\chi\wedge\phi)\vee(\neg\chi\wedge\phi)$
4     $\leftrightarrow(\chi\vee\neg\chi)\wedge\phi$ Distributivity
5     $\leftrightarrow\phi$ Theorem (LEM) like top
6 $(\chi\wedge\phi)\vee(\neg\chi\wedge\phi)\leftrightarrow\phi$ 4–5: Transitivity
7 $\phi\to\psi$ 6,3: Replacement
Rule 3.4.19 (Resolution).
$\{\phi\vee\chi,\psi\vee\neg\chi\}\Rightarrow\phi\vee\psi.$
@@ 1 $\neg\chi\wedge(\phi\vee\chi)\to\phi$ Disjunctive syllogism
2 $\chi\wedge(\psi\vee\neg\chi)\to\psi$ Disjunctive syllogism
3 $\neg\chi\wedge(\phi\vee\chi)\to\phi\vee\psi$ 1: Strengthening
4 $\chi\wedge(\psi\vee\neg\chi)\to\phi\vee\psi$ 2: Strengthening
5 $\neg\chi\wedge(\phi\vee\chi)\wedge(\psi\vee\neg\chi)\to\phi\vee\psi$ 3: Weakening
6 $\chi\wedge(\phi\vee\chi)\wedge(\psi\vee\neg\chi)\to\phi\vee\psi$ 4: Weakening
7 $(\phi\vee\chi)\wedge(\psi\vee\neg\chi)\to\phi\vee\psi$ 5,6: LEM exhaustion for implication
Rule 3.4.20 (Cut).
$\{\phi_{1}\to\chi,\phi_{2}\wedge\chi\to\psi\}\Rightarrow\phi_{1}\wedge\phi_{2}\to\psi.$